Two Jets are traveling torward each otherand are 4000km apart. The rate of one jet is 100 km/h faster than the rate of the other. If the jets pass each other after 2.5 hours, what is the rate of the faster jet
Answer by checkley71(8403)

SORRY DIDN'T READ THE ENTIRE PROBLEM. THEY ARE ASKING FOR THE FASTEST WHICH WOULD BE 750=100 OR 850 KM/H.
DISTANCE = RATE TIMES TIME
4000=[X+(X+100)]2.5
4000=(2X+100)2.5
4000=5X+250
5X=4000-250
5X=3750
X=3750/5
X=750
PROOF
4000=(750+850)2.5
4000=1600*2.5
4000=4000